OpenStudy (anonymous):

A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). There are two production methods it could use. With one method, the one-time fixed costs will total $14,532 , and the variable costs will be $19.25 per book. With the other method, the one-time fixed costs will total $34,447 , and the variable costs will be $10.50 per book. For how many books produced will the costs from the two methods be the same?

OpenStudy (seascorpion1):

you should graph these as a y=ax+b (slope-intercept form): \[y_1=14532+19.25b\] \[y_2=34447+10.5b\] where b is the number of books. Then set \[y_1=y_2\] and solve for b.